OpenStudy (logan13):

Aidan bought the roll of paper shown below for a school project. The radius of the roll is 7 centimeters. Which expression can be used to find the approximate circumference of the roll of paper?

OpenStudy (logan13):

https://app159.studyisland.com/pics/216030n1.png

OpenStudy (logan13):

@Unanimoose

OpenStudy (logan13):

want to help

OpenStudy (logan13):

can you fan me uniamoose

OpenStudy (unanimoose):

\[A = D \pi\]

OpenStudy (unanimoose):

Where D is the diameter. The radius doubled is the diameter.

OpenStudy (unanimoose):

So: \[A = (7 \times 2)\pi\]

OpenStudy (logan13):

so a

OpenStudy (logan13):

@Unanimoose

OpenStudy (unanimoose):

I can't see the options.